Solve by completing the square x^2+6x=12
Show work please, thanks

Answer by stanbon(75887)

Solve by completing the square x^2+6x=12
-----
x^2 + 6x + 9 = 12+9
-----
(x+3)^2 = 21
--------
x+3 = sqrt(21) or x+3 = -sqrt(21)
------
x = -3+sqrt(21) or x = -3-sqrt(21)
----------------
Cheers,
Stan H.
----------------